### v1

# Periodic Chains Scheduling on Dedicated Resources - A Crucial Problem in Time-Sensitive Networks

Abstract
Periodic messages transfer data from sensors to actuators in cars, planes, and complex production machines. When considering a given routing, the unicast message starts at its source and goes over several dedicated resources to reach its destination. Such unicast message can be represented as a chain of point-to-point communications. Thus, the scheduling of the periodic chains is a principal problem in time-triggered Ethernet, like IEEE 802.1Qbv Time-Sensitive Networks. This paper studies a strongly NP-hard periodic scheduling problem with harmonic periods, task chains, and dedicated resources. We analyze the problem on several levels of generality and complexity and provide the corresponding proofs. We describe a solution methodology to find a feasible schedule that minimizes the chains’ degeneracy related to start-to-end latency normalized in the number of periods. We use the local search with the first fit scheduling heuristic, which we warm-start with a constraint programming model. This notably improves the schedulability of instances with up to 100% utilization and thousands (and more) of tasks, with high-quality solutions found in minutes. An efficient constraint programming matheuristic significantly reduces the degeneracy of the found schedules even further. The method is evaluated on sets of industrial-, avionic-, and automotive-inspired instances.
keywords:
periodic scheduling , constraint programming , bin packing , Time-Sensitive Networks
journal: Computers & Operations Research
[…]
{highlights}
Proof of Strong NP-hardness
Constraint Programming Warm Start
Constraint Programming Matheuristic
1 Introduction
1.1 Motivation
Periodic events and processes play an essential role in many application domains. To ensure efficient and deterministic data exchange among different parts of the application, engineers need to configure communication protocols to satisfy timing and throughput requirements. For example, with the advent of self-driving cars [1] requiring a deterministic exchange of high-volume data, the communication network may quickly become a congestion point, and efficient schedules are needed to satisfy all safety-relevant features. When considering a given routing, the periodic (unicast) message goes from a sensor over several dedicated resources to reach its actuator. Thus, the message can be represented as a chain of point-to-point communications performed on communication links.
Ethernet is widely used in everyday life due to its high throughput and availability of SW stacks. Still, it lacks time determinism, essential for real-time applications, such as in-vehicle networks, industrial production lines, and audio streaming, that are very sensitive to delays. Time-sensitive networking (TSN) is a set of new standards that enables sending messages in a time-triggered manner. To use TSN efficiently and satisfy various requirements, many researchers [2, 3] have developed algorithms that can synthesize a schedule of periodic messages. Their results are versatile, but the network utilization and algorithm performance are often very low. Thus, in this paper, we omit a complex behavior of the TSN (queues, deadlines, frame isolation), and we concentrate on a crucial part of this problem: the scheduling of periodic tasks. We focus on harmonic periods of the messages (i.e., for each pair of distinct periods, it holds that the longer period is divisible by the shorter one) that are widely used in engineering practice and help with the tractability of the Periodic Scheduling Problem (PSP).
1.2 Related Work
Scheduling of periodic tasks has been studied from many different directions, as is outlined in [4]. Crucial results regarding the complexity of the scheduling of periodic tasks on a single resource were presented in [5], where the authors proved that the problem with non-preemptive tasks is strongly NP-complete. However, in [6], it was shown that if both the periods and processing times of the tasks are harmonic, a feasible schedule can be found in polynomial time.
Scheduling of both harmonic and non-harmonic periodic problems, as well as some special cases, was studied in [7, 8]. The authors presented a local schedulability condition for the general case of PSP. Motivated by the desire to increase the schedulability for highly-utilized instances, alternative views of the PSP based on bin-trees [9] and bin packing [10, 11, 12] were presented. These perspectives enabled the relevant authors to devise heuristics and exact methods to efficiently find feasible schedules.
The PSPs are often found in the domains of avionics and automotive. Building on their previous work, authors of [13] used Integer Linear Programming (ILP) to partition tasks on different processors of modern aircraft. In [14], the goal was to schedule tasks in the avionics domain to maximize a leeway between them. This way, the occasional increase in the processing time of the task would not interfere with the rest of the schedule. A game-theoretic approach to tackle a principally similar problem was presented in [15]. In [16], the authors tackled the scheduling of so-called partitions on shared processors. Their idea utilized a greedy heuristic, streamlining an otherwise exhaustive approach without a significant loss in the quality of the schedules. In [17], communication between modules of the aircraft is considered, and the authors solve it with constraint-generating ILP procedure.
In automotive domains, many papers tackle scheduling problems related to specific protocols and architectures. In [18], the authors schedule tasks for AUTOSAR architecture to eliminate load peaks in time. ILP and heuristics for the same architecture were also investigated in [19]. FlexRay, a protocol often coupled with the AUTOSAR specification, was studied in [10] and more recently in [20, 21]. In the latter, the authors consider multiple message variants so the same schedule can be repurposed for different vehicle models.
In the previous paragraphs, the authors mostly focused on partitioning tasks on different resources and finding appropriate schedules. However, in cases with messages passing from one resource to another, the assignment of the tasks to the resources is determined beforehand during the routing phase.
In [22], the authors co-optimize the resource utilization as well as the control performance of the system that relies on signals from sensors passed through the network. In [23], large-neighborhood search is used to improve control-wise properties of the schedules.
Early work on offline scheduling for TTEthernet (an example of time-triggered ethernet) was presented in [24]. The authors studied several network topologies, including line topology, and used satisfiability modulo theories (SMT) solver in their proposed incremental workflow. Many papers emerged with the introduction of the TSN; see survey papers [2, 3]. In the latter, the authors categorized several different methods and performed experimental comparisons.
In [25], the authors developed an SMT model to shorten gate control lists controlling the schedules; their length is a limited resource in practice. Characteristics of the real industrial networks were exploited in [26] to develop a two-stage scheduling method; authors were able to isolate parts of the problem and solve them independently of each other. The TSN formalism can be further extended to include, e.g., authentication, as is demonstrated in [27]. Cryptography keys are disclosed with a time delay, enabled by the TSN, so nodes of the network may authenticate themselves. In our previous TSN papers [28, 29], concentrating on a complete set of TSN requirements and tight time windows, we achieved very low utilization of the communication resources.
On the other hand, in [30], where we presented a first-fit heuristic for a very similar problem (dedicated resources with soft constraints in the form of criteria instead of hard ones), we achieved nearly 100% utilization of the resource. This observation triggered the research presented in this paper, concentrating on the principal aspects of scheduling periodic chains on dedicated resources.
Apart from communications, PSPs apply to manufacturing and maintenance. In [31, 32], the scheduling problems with (almost) periodic maintenance windows are studied; there, the maintenance task is frequently repeated, while the manufacturing tasks are essentially aperiodic. This framework is useful in practice since inspecting the machines periodically reduces the risk of failures. A practical application can be found in [33], where the authors schedule patients’ non-preemptive treatments around the medical machine’s maintenance.
1.3 Outline and Contributions
In this work, we focus on the scheduling of messages, represented by chains of non-preemptive tasks. Each task is to be performed on a dedicated resource, and all tasks of the chain have the same period from a set of harmonic periods. We primarily optimize the degeneracy, which is related to the message’s start-to-end latency normalized in the number of periods. We build upon our single resource scheduling paper [12] and extend our heuristic outlined in [30].
In Section 2, we formally describe the problem and its structural properties, including the connection between PSP with harmonic periods and bin-packing. In Section 3, we discuss the complexity of several special cases of PSPs. We prove it is polynomial for a single period, whereas it is strongly NP-Hard for two periods. In Section 4, we describe a heuristic algorithm and Constraint Programming (CP) approaches to warm-start the algorithm and fine-tune the final solution. In Section 5, we discuss the experimental results, and in Section 6, the conclusions are drawn.
The main contributions of our paper are:
  • 1.
    Proofs of the complexity of several classes of PSPs
  • 2.
    CP warm-start approach for highly-utilized instances
  • 3.
    CP-based matheuristic improving found schedule
2 Problem Description
2.1 Problem Statement
Tasks and Resources
We consider a set of $\kappa$ chains. A chain $\mathcal{C}^{k}$ with $k\in\{1,\ldots,\kappa\}$ is composed of tasks $J_{i}^{k}$ $i\in\{1,\ldots,n^{k}\}$ with integer processing times $p_{i}^{k}$. Each task is to be performed on a pre-defined unit capacity resource among $\{1,\ldots,M\}$. The index of the resource processing $J_{i}^{k}$ is denoted by $\pi_{i}^{k}$.
In TSN applications, a chain corresponds to a message sent periodically, and resources are successive communication links used to deliver the message. A point-to-point communication performed by such communication link corresponds to a task. We elaborate on this correspondence in Fig. 1(a). Notice that since the resource corresponds to a communication link rather than a node, the messages in the example consist of three point-to-point communications (i.e., tasks) but travel between four nodes of the network. From a scheduling point of view, a resource can be viewed as a machine. Our problem thus shares some features with job-shop problems.
The tasks of a chain are periodic and share the same period. We assume that the periods belong to a harmonic set $\tau(1),\ldots,\tau(\rho)$ of size $\rho$, so that $\forall l\,\in\nolinebreak\{2,\ldots,\rho\},\tau(l)=b_{l}\tau(l-1)$ where $b_{l}$ is an integer. We define the least period $w=\tau(1)$, and we define the hyperperiod $\hat{H}$ to be the least common multiple of periods. Due to the harmonic nature of the period set, $\hat{H}$ is equal to the longest period $\tau(\rho)$. All of the tasks of chain $\mathcal{C}^{k}$ have period $T^{k}=\tau(\beta^{k})$, where $\beta^{k}$ is the index of period $T^{k}$ in the ordered set of periods.
For resource $m$ we denote by $\mathcal{J}_{m}$ the set of tasks $J_{i}^{k}$ assigned to it (i.e. such that $\pi_{i}^{k}=m$). We define the utilization of resource $m\in\{1,\ldots,M\}$ as:
$U_{m}=\sum_{J_{i}^{k}\in\mathcal{J}_{m}}\frac{p_{i}^{k}}{T^{k}}$
The utilization $U_{m}\leq 1$ represents the percentage of time resource $m$ is busy in any feasible schedule. The utilization of an instance is the maximum utilization of its resources.
Notation
According to the standard Graham notation extended to handle PSPs in [4], our problem is denoted by ${PD|T_{i}^{harm},jit=0,chains|criteria}$. $PD$ stands for parallel dedicated processors, $chains$ for chains, $T_{i}^{harm}$ for harmonic periods, $jit=0$ since no jitter is allowed (i.e., the starting time of a task is entirely determined by the starting time of its first occurrence and by its period). We assume that all tasks of a chain have the same period.
(a) Ten messages are sent from the leftmost node A to the rightmost node D via three communication links (i.e., resources). Corresponding chains $\mathcal{C}^{1}-\mathcal{C}^{10}$ consisting of three tasks $J_{i}^{k}$ are to be scheduled on resources 1, 2, 3 (assignment of tasks to resources $\pi_{i}^{k}=i$ is shown by dotted ellipses). For simplicity, no messages are sent via links in the opposite direction 4, 5, 6.
(b) Schedule of 10 chains on 3 resources spanning over one hyperperiod. Tasks of the chain have the same color. Total degeneracy is 2. Start-to-end latency $S^{k}$ is shown for cyan $\mathcal{C}^{1}$ and yellow $\mathcal{C}^{3}$ chains, which both have degeneracy $D^{k}=1$.
Figure 1: Illustrative communication case and associated schedules for instance with periods $\tau(1)=40,\tau(2)=80,\tau(3)=240$. 10 chains are to be scheduled; $\mathcal{C}^{1}-\mathcal{C}^{2}$ have period 40 (thus $T^{1}=T^{2}=\tau(1)$, meaning $\beta^{1}=\beta^{2}=1$), $\mathcal{C}^{3}-\mathcal{C}^{7}$ have period 80, and $\mathcal{C}^{8}-\mathcal{C}^{10}$ have period 240. Notice that all the chains are sent from left to right in this example; communication links in the other direction are not used. Utilization $U_{m}=0.96$ holds for all resources.
Schedules
A schedule assigns an integer start time $t_{i}^{k}$ to the first occurrence of each task $J_{i}^{k}$, such that the precedences are fulfilled:
$t_{i+1}^{k}\geq t_{i}^{k}+p_{i}^{k}\quad\forall k\in\{1,\ldots,\kappa\},\forall i\in\{1,\ldots,n^{k}-1\},\quad$ (1)
and at any time instant $s$, a processor performs at most one task:
|$\left|\left\{J_{i}^{k}\in\mathcal{J}_{m},\quad\exists l\in\mathbb{N},\quad t_{i}^{k}+l\cdot T^{k}\leq s<t_{i}^{k}+p_{i}^{k}+l\cdot T^{k}\right\}\right|\leq 1\\ \forall s\geq 0,\forall m\in\{1,\ldots,M\},$ (2)
Notice that if independent periodic tasks $\mathcal{J}_{m}$ can be scheduled on their respective resources, we can derive a schedule that satisfies precedence constraints as well. Indeed, assume we are given a feasible schedule $\sigma$ of the independent periodic tasks where $\sigma_{i}^{k}<T^{k}$ is the start time of the first occurrence of $J_{i}^{k}\in\nolinebreak\mathcal{J}_{m}$. Then, we can satisfy the precedence constraints (1) of each chain $k$ by postponing a subsequent task in the chain to the next period whenever needed to satisfy the precedence constraint. Formally, we can set:
$t_{i}^{k}=\sigma_{i}^{k}+q_{i}T^{k}$
with $q_{1}=0$. If $\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}$, then we set $q_{i+1}=q_{i}$. If $\sigma_{i+1}^{k}<\sigma_{i}^{k}+p_{i}^{k}\leq\sigma_{i+1}^{k}+T^{k}$ we set $q_{i+1}=q_{i}+1$. Otherwise we set $q_{i+1}=q_{i}+2$. Observe that in the latter case, we know that $\sigma_{i}^{k}+p_{i}^{k}<2\cdot T^{k}\leq\sigma_{i+1}^{k}+2\cdot T^{k}$. So, in any case, $t_{i}^{k}+p_{i}^{k}\leq t_{i+1}^{k}$ so that the precedence constraints are satisfied.
Conversely, a feasible integer schedule of the chains induces a feasible schedule $\sigma$ of independent periodic tasks of $\mathcal{J}_{m}$ on resource ${m}$ by setting
$\sigma_{i}^{k}=t_{i}^{k}\mod T^{k}.$
with the additional property that $\sigma_{i}^{k}<T^{k}$. This schedule is called the core schedule of resource $m$.
Such schedules, shown within a single hyperperiod, can be illustrated as in Fig. 1. There, 10 chains are delivered using a sequence of resources 1, 2, and 3. Precedence constraints for chains hold, and no overlaps occur on either resource.
Criteria
All criteria we consider are related to the start-to-end latency of chains, i.e., the delay between the moment when the chain’s first task started and the chain’s last task ended. In communications, this refers to the delay between the moment the message was sent from the first (source) resource and when the message was processed at the last (sink) resource.
Let $t$ be a schedule. We define the following criteria:
Start-to-end latency of chain $k$:
$S^{k}=t_{n^{k}}^{k}+p_{n^{k}}^{k}-t_{1}^{k}$
Degeneracy of chain $k$:
$D^{k}=\left\lceil\frac{S^{k}}{T^{k}}\right\rceil-1$. The degeneracy is the start-to-end latency normalized in the number of periods minus 1. Thus, a chain with start-to-end latency shorter than its period has degeneracy zero.
$\alpha$-degeneracy of a chain $k$:
Given an $\alpha\in(0,1]$ $D^{k}(\alpha)=\left\lceil\frac{S^{k}}{\alpha\cdot T^{k}}\right\rceil-1$. This criterion uses finer granularity of the degeneracy whose minimization might be more restrictive: for example, $D^{k}(0.5)=0$, which means that the start-to-end latency of chain $\mathcal{C}^{k}$ fits in half of the chain period.
Notice that the degeneracy of a chain can be computed by considering the core schedule of the chain and the number of postponing operations necessary to satisfy the precedence constraints. So, the degeneracy of the chain is the value of $q_{n^{k}}$ mentioned earlier.
For each measure, we consider the max value and the sum of values:
  • 1.
    max degeneracy $D_{\max}=\max_{k=1,\ldots,\kappa}D^{k}$ and degeneracy $D_{sum}=\sum_{k=1}^{\kappa}D^{k}$.
  • 2.
    similarly max and sum $\alpha$-degeneracy can be defined $D_{\max}({\alpha}),D_{sum}({\alpha})$.
2.2 Special Cases
In the paper, we consider both general $PD|T_{i}^{harm},jit=0,chains|criteria$ instances, as well as instances generated with a set of realistic assumptions (e.g., using underlying network topology) to fit the practical applications where chains represent messages sent over communication links. Some of these assumptions, which will be used in the following sections of the paper, are:
Limited number of resources
This is described in the first field of Graham notation. For example, $PD2$ indicates two resources.
Harmonic processing times.
In this case, all values of processing times and periods are together in a harmonic set. This makes the single resource feasibility problem polynomial [6]. Graham notation for such case is $PD|(T_{i},p_{i})^{harm},jit=0,chains|criteria$.
Topology
In this case, the communications links are organized along a topology (line, tree, for example). A chain is characterized by the source resource and the sink resource, and the chain’s tasks are assigned to the sequence of resources on the unique path from the source to the sink.
Equal processing time on a chain
In this case, inspired by the most common case in communications, point-to-point communications have the same duration since the packets have the same length and communication links have the same speed. In such case $p_{i}^{k}=p^{k}$.
Scattering
A special case where all the first tasks of the chains are assigned to the same resource; we call this scattering (as in the context of parallel computing [34]). Proofs related to this case are described in Section 3. Figure 1 shows a scattering example over a line topology with equal processing times.
2.3 Structural Properties
In [12], we established a formal relation between periodic schedules on a single resource and a particular case of 2D packing problem. The packing view is a crucial inspiration behind the warm-start approach described later.
2.3.1 Single Resource Core Schedule Coordinates
Consider a single resource $m$ and a feasible core schedule $\sigma$ of the independent periodic tasks of $\mathcal{J}_{m}$. Let $o_{m}$ be the start time of the first task $J_{f}^{k}$ on resource $m$ with the least period $w$. We consider $o_{m}$ as the origin for resource $m$. So we define the shifted core schedule $s_{i}^{k}$ for each task $J_{i}^{k}\in\mathcal{J}_{m}$ such that:
$s_{i}^{k}=\left\{\begin{array}[]{ll}\sigma_{i}^{k}-o_{m}&if\ \sigma_{i}^{k}\geq o_{m}\\ \sigma_{i}^{k}+T^{k}-o_{m}&otherwise\end{array}\right.$
We call time interval $[o_{m},o_{m}+\hat{H})$ the ”observation interval.” Notice that for any integer $v\geq 1$, the schedule $\sigma$ of time interval $[o_{m}+(v-1)\hat{H},o_{m}+v\hat{H})$ on resource $m$ is the same as the schedule of the observation interval shifted by $(v-1)\hat{H}$ time units. This offset may differ for different resources.
2.3.2 Associated HD2D Packing Problem
We can now define a height-divisible 2D (HD2D) packing instance associated with a periodic scheduling instance with harmonic periods. Consider an instance $I$ of the scheduling problem.
To each task $J_{i}^{k}\in\mathcal{J}_{m}$ we associate a rectangle $R_{i}^{k}$ of width $\ell_{i}^{k}=p_{i}^{k}$ and height $0pt{i}{k}=\frac{\hat{H}}{T^{k}}$, to fit in a bin of width $w$ and height $H=\frac{\hat{H}}{w}$ without overlapping rectangles. Moreover, the packing has to be height divisible, which means that the vertical coordinate of each rectangle is a multiple of its height.
(a) Schedule of a single observation interval of resource 2 arranged in rows of length 40
(b) Height-divisible 2D packing of resource 2
Figure 2: Observation interval and packing perspective of a schedule of resource 2 from Fig. 1.
Paper [12] gives the formal correspondence between the start time of a task in the core schedule $s_{i}^{k}$ and coordinates of the bottom left corner of the associated rectangle. To give an intuition of this correspondence, we develop it on an example in Fig. 2. The main idea is to arrange the periodic schedule in rows. A row indexed by $v$ represents the schedule of interval $[vw,(v+1)w)$ of the observation interval.
In Fig. 2(a), the single ”observation interval” associated with the schedule of resource 2 from motivational example Fig. 1 is shown. Ten chains $\mathcal{C}^{1},\dots,\mathcal{C}^{10}$ pass through the resource. $\mathcal{C}^{1},\mathcal{C}^{2}$ have the shortest period $\tau(1)=40$. Chains $\mathcal{C}^{3}-\mathcal{C}^{7}$ have a period equal to 80; the rest have a period equal to 240, which is the hyperperiod. Origin 00 is set by the start time of chain $\mathcal{C}^{1}$, and we stack each interval of length $w=40$ on top of the previous one. So row $v$ describes the schedule of interval $[v\cdot w,(v+1)w)$. Horizontal coordinate $u$ corresponds to the relative position of the task’s occurrence in the interval.
Row indices are represented by two-digit numbers using a mixed-radix system based on the period multipliers 2 and 3: here period $\tau(2)=2\cdot\tau(1)$, and $\tau(3)=3\cdot\tau(2)$.
Figure 2(b) shows the corresponding height-divisible packing view of the same solution. Coordinates of Rectangles are described with coordinates of their bottom-left corners $(x,y)$. The packing is obtained by interchanging the rows as follows: we swap the two-digit row numbers in Figure 2(a) such that the most significant digit becomes the least significant digit and vice versa. Finally, we obtain Figure 2(b) by ordering the rows according to their new two-digit numbers. For example, the row with $\mathcal{C}^{6}$ has number $10_{3}1_{2}=$ 1 in Figure 2(a) and after the swap, the same row has number $31_{2}0_{3}=$ 3 in Figure 2(b). We can see the occurrences of the original tasks have been merged, and the dimensions of the resulting rectangles correspond to the rectangles of the associated packing instance. For example, tasks 1-2 with the shortest period are transformed into rectangles spanning the entire height of the bin. Tasks 3-6 merge three of their occurrences into a rectangle with height 3, and finally, rectangles associated with the longest-period tasks 8-10 have a height equal to 1.
Note that thanks to the equivalence between the two problems, we can recover a periodic schedule by providing a solution for the height-divisible packing problem and vice versa.
3 Complexity Analysis
In this section, we consider related complexity results to explain the complexity of our problems. First, we know that scheduling periodic tasks with harmonic periods on a single resource is a strongly NP-hard problem [5]. So, minimizing one of our criteria seems at least as difficult as this problem and is still NP-hard unless very strong assumptions are made. In the rest of the section, we show some very special polynomial cases of our problem and present a strongly NP-hard subproblem.
3.1 Scattering on a Line Topology with Two Resources
Let us consider a special case of the line topology with two resources and a scattering hypothesis. So, all chains have their first task on resource 1 and second on resource 2.
3.1.1 Single Period
Assume all chains have the same period $T$. Then, finding a feasible schedule for a single resource becomes an easy problem. We can prove that as for the two-machine flow-shop problem, minimizing the degeneracy can be done in polynomial time. Notice that in this case, the sum of processing times of tasks on any resource is not greater than $T$. Otherwise, the problem would be infeasible.
Lemma 1.
Scattering on a line of two resources for $PD2|T_{i}=T,jit=0,chains|D_{sum}$ and $PD2|T_{i}=T,jit=0,chains|D_{\max}$ can be done in polynomial time.
Proof.
The non-periodic version of this problem is a two-machine flow-shop. A chain $\mathcal{C}^{k}$ corresponds to a job with two operations $J_{1}^{k},J_{2}^{k}$ performed on the first and second machine, respectively. And we know from [35], a schedule that optimizes the makespan ($C_{max}=\max_{k\in\{1,\ldots,\kappa\}}t_{2}^{k}+p_{2}^{k}$) can be computed in polynomial time using Johnson’s rule. Moreover, we know that it is a permutation schedule: the order of jobs is the same on both machines even though it was not required. If we shift the schedule of the first machine to the left and of the second machine to the right so that there does not remain idle time between two tasks on the resources and it ends at the optimal makespan, then the schedule can be repeated every $T$ time units without conflicts. We prove that this minimizes degeneracy.
Indeed, consider a pivot chain $\mathcal{C}^{k^{*}}$, for which $t_{1}^{k^{*}}+p_{1}^{k^{*}}=t_{2}^{k^{*}}$. Such a pivot chain exists. Otherwise, the schedule on the second machine could be left shifted, and thus, the makespan would not be optimal. The start-to-end latency for chain $\mathcal{C}^{k^{*}}$ is thus minimal, and so is its degeneracy. Then for any chain $k$ performed before this $k^{*}$ in the schedule, necessarily the start-to-end latency $S^{k}\leq t_{2}^{k^{*}}$ which is not greater than the sum of processing times on the first resource, which is not greater than $T$. Hence $D^{k}=0$. Symmetrically, for a chain $l$ performed after chain $k^{*}$ the value of $S^{l}$ should be less than the sum of processing times on the second resource, which is not greater than $T$ so that $D^{l}=0$. This proves that this schedule minimizes both $D_{\max},D_{sum}$. ∎
3.1.2 Two Periods
As in Section 3.1.1, we assume scattering chains on two resources, but now the chains can have two different harmonic periods. In [6], the authors give a reduction from $3-PARTITION$ to a single resource scheduling problem with two different harmonic periods, thus proving that the feasibility of a single resource PSP with two periods is strongly NP-Hard.
From this result, we can deduce the following:
Property 1.
For scattering chains on two resources in line topology, finding a schedule with $D_{sum}$ or $D_{\max}$ less than a value $d$ is strongly NP-hard.
Proof.
Consider an instance of the single resource scheduling problem with two harmonic periods. Period $T^{i}$ of task $i$ is equal to one of the harmonic periods $\tau(1)$ or $\tau(2)$. Notice that the processing time of $i$ can be assumed strictly less than its period; otherwise $i$ would be the only task scheduled on the resource. For each task $i$ of the single resource problem, we define a chain $\mathcal{C}^{i}$ with period $T^{i}$ composed of task $i$ on the first resource and a task with processing time $1$ on the second resource. Obviously, if a feasible schedule for the single resource instance exists, it gives a schedule for the two resources of the chains with 00 degeneracy by scheduling the tasks on the second resource as soon as possible with respect to precedence constraint. Conversely, a schedule with 00 degeneracy implies the existence of a feasible schedule on the first resource. So this defines a reduction from the single resource, two harmonic periods problem to the scattering problem on two resources with two periods. ∎
As was shown in [6], the single resource problem with processing times and the periods belonging to a harmonic set is polynomial. However, the complexity of the two-resource scattering on a line case where periods and processing times are from one harmonic set remains open.
3.2 Scattering on a Line or an Out-Tree
In this section, we assume equal processing times of all tasks of a chain, and that the resources are organized along a given topology. We first consider the special case of scattering along a line.
3.2.1 Scattering on a Line
In this subsection, we consider a line of resources indexed from $1$ to $M$, and we suppose that each chain $\mathcal{C}^{k}$ starts at resource $1$ and ends at resource $u^{k}$. The chain comprises of tasks $J_{i}^{k}$ with $1\leq i\leq u^{k}$. Let us denote by $p^{k}$ the processing time of every task of chain $\mathcal{C}^{k}$.
So resource $1$ processes all the tasks from all chains $\mathcal{C}^{k},k=1,\ldots,\kappa$. Then, if there exists a feasible schedule for resource $1$, there exists a feasible schedule for each resource and thus a feasible schedule for the problem.
We call an offset schedule a periodic schedule built by assuming that for any resource $i$ a chain $\mathcal{C}^{k}$, task $J_{i}^{k}$ with $i\leq u^{k}$ is scheduled exactly as it is in resource $1$, with an offset that only depends on $i$. We define $o_{i}$ to be the offset between schedules of resource $1$ and resource $i$. Then, for each chain $\mathcal{C}^{k}$ with $u^{k}\geq i>1$ it holds that $t_{i}^{k}=t_{1}^{k}+o_{i}$.
Property 2.
Suppose we are given a feasible schedule $\sigma$ of the tasks on the first resource. The minimal degeneracy $D_{\max}$ or $D_{sum}$ of an offset schedule $\omega$ associated with $\sigma$ can be computed in polynomial time. In this schedule, we have the following inequality:
$D_{\max}(\omega)\leq\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1)$
where $p^{\min}=\min_{k}p^{k},p^{\max}=\max_{k}p^{k}$ and $D_{\max}(opt)$ is the minimal maximum degeneracy.
Proof.
Let $i>1$ be a resource. We can prove that in any feasible offset schedule, $o_{i}-o_{i-1}\geq\max_{k,u^{k}\geq i}p^{k}$. Thus $o_{i}\geq i\cdot\max_{k,u^{k}\geq i}p^{k}$. Indeed, let $\mathcal{C}^{k^{*}}$ be the chain with maximum processing time among chains with $u^{k}\geq i$. We know that $J_{i-1}^{k^{*}}$ precedes $J_{i}^{k^{*}}$ so that $t_{i}^{k}\geq t_{i-1}^{k}+p^{k^{*}}$. Notice that in any schedule, the start-to-end latency of a chain $\mathcal{C}^{k}$ equals $S^{k}=o_{u^{k}}+p^{k}$, So in any offset schedule, the minimum start-to-end latency of the chains is obtained by stating $o_{i}=i\cdot\max_{k,u^{k}\geq i}p^{k}$ for each $i$.
Consider now any chain $\mathcal{C}^{k}$. In any schedule $s$, $S^{k}(s)\geq n^{k}p^{k}\geq n^{k}p^{\min}$. And with the proposed offset schedule $\omega$, $S^{k}(\omega)=(n^{k}-1)p^{\max}+p^{k}\leq n^{k}p^{\max}$. So, $\frac{S^{k}(\omega)}{S^{k}(s)}\leq\frac{p^{\max}}{p^{\min}}$.
Now, for some chain $\mathcal{C}^{k}$,
$D_{\max}(\omega)=\left\lceil\frac{S^{k}(\omega)}{T^{k}}\right\rceil-1\leq\frac{S^{k}(\omega)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}\frac{S^{k}(opt)}{T^{k}}\leq\frac{p^{\max}}{p^{\min}}(D_{\max}(opt)+1)$
If we assume harmonic periods and processing times (making the single resource problem polynomial), Lemma 2 gives a simple polynomial algorithm to compute a good offset schedule for criteria $D_{sum}$ and provides a performance guarantee for criteria $D_{\max}$.
When processing times are arbitrary, we can still apply the offset schedule as long as we are able to find a feasible schedule using, e.g., CP or SMT. First, we find a feasible schedule for the first resource of the line, and then we create schedules for other resources by applying a suitable offset $o_{i}$. Even though the single resource problem remains strongly NP-hard, we empirically show in Section 5.3 that the offset schedule combined with the CP model for the single resource problem works very well.
3.2.2 Periodic Scattering along a Tree
Similar offset schedules can be used on scattering on an out-tree topology, assuming that the first task of each chain is performed on the root resource of the tree and the following tasks use consecutive resources along a path of the tree. This is shown in Fig. 3. Notice that there are actually two root resources, A-B and A-F, which lead to a decomposition of the instance to scattering along a tree (chains $\mathcal{C}^{1},\mathcal{C}^{2}$) and along a line (chain $\mathcal{C}^{3}$). Only the first tasks of the chains $\mathcal{C}^{1},\mathcal{C}^{2}$ share the same resource A-B.
Figure 3: Example of a scattering along a tree instance. Chains $\mathcal{C}^{1},\mathcal{C}^{2}$ originate on link A-B, and their paths diverge afterward. Chain $\mathcal{C}^{3}$ is completely independent since it originates on a different link A-F.
Similarly, as in the previous section, we can duplicate the schedule of the root on the intermediate nodes, removing the chains that are not processed and defining offsets with respect to the schedule of the root resource. Similarly, if resource $j$ is a successor of resource $i$ in the tree, then for a chain $\mathcal{C}^{k}$ that uses this communication link, we should have $o_{j}-o_{i}\geq p_{i}^{k}=p^{k}$. So that the optimal degeneracy of offset schedules based on a given schedule of the root resource is easy to compute. Analogously, this approach can be used for gathering information in the sink of an in-tree.
4 Solution Approach and Algorithms
Figure 4: Outline of the proposed approach predecessor-flow.
This section describes the proposed solution approach for solving PSPs with harmonic periods. The approach is based on our local search [30], which uses the first fit scheduling (FFS) heuristic. The overall flow, visualized in Fig. 4, shows the local search and our two CP-based improvements. If the local search (Section 4.1) is unable to find a feasible solution quickly, the first improvement provides it with the warm start (Section 4.2); this is particularly useful for solving highly-utilized instances. The second improvement reduces the degeneracy of the found schedule even further using a matheuristic (Section 4.3).
4.1 Local Search
We outlined the local search method in [30]. The inner FFS heuristic processes the tasks following the ordered list of tasks $\mathcal{J}$; the ordered lists are generated and supplied to the heuristic by the encompassing local search. One by one, the tasks are scheduled on their dedicated resources as shown in Algorithm 1.
We distinguish two methods. The leftmost method schedules each task as early as possible so that none of its occurrences overlap with any occurrence of any already scheduled tasks (for more details, see Section 4 in [30]). This method is more suitable for finding a feasible solution, but it may result in a large degeneracy since it ignores the precedence constraint. The second predecessor method performs the following: If predecessor $i$ of the task $i+1$ in chain $k$ is already scheduled, we schedule the current task only after the completion time of its predecessor. Unlike the previous method, the predecessor method does not ignore the precedence of already scheduled predecessors. In both cases, the actual schedule is obtained by postponing the tasks that do not adhere to the precedence constraints.
1
Input : tasks $J_{i}^{k}$ with periods, processing times, and assignment to resources; chains $\mathcal{C}^{k}$ with precedences; method; ordered list $\mathcal{J}$
Output : schedule; or infeasible $\mathcal{J}$
while $\mathcal{J}\neq\{\}$ do
       $J_{i}^{k}\leftarrow\mathcal{J}$.pop() // remove first task from the ordered list
       if method $=$ predecessor then
             $J_{i-1}^{k}\leftarrow$ predecessor of $J_{i}^{k}$
            if $J_{i-1}^{k}$ exists and already scheduled then
                   $t\leftarrow t_{i-1}^{k}+p_{i-1}^{k}$
            else
                  $t\leftarrow 0$
             end if
      else
            $t\leftarrow 0$ // method = leftmost
       end if
      schedule $J_{i}^{k}$ as early as possible starting from $t$
      if $J_{i}^{k}$ is not scheduled then
            return infeasible $\mathcal{J}$
       end if
end while
postpone tasks to satisfy precedence constraints
return schedule
Algorithm 1 First fit scheduling heuristic
We use local search to find the appropriate ordered list $\mathcal{J}$ to provide to the FFS heuristic. Starting with an initial ordered list, the local search first tries to modify it to respect the precedences of chains. Once this fails, the local search performs either random moves or chain reordering. The random moves are performed by swapping two tasks in the ordered list. The reordering operation takes one (arbitrary) chain whose order of tasks in the list does not follow its precedence constraints and reorders these tasks to follow them.
In the first iteration, the local search begins with a rate-monotonic ordered list, which sorts tasks in ascending order of their periods. As we discussed in [12], the rate-monotonic ordered list is the desirable starting point since we process most constrained tasks (i.e., tasks with most occurrences) at the start of the procedure when the schedule is mostly empty.
4.2 Warm-Start by Constraint Programming
The local search approach generally works very well. However, with the utilization of instances close to 100%, the approach struggles to find any feasible solution. In that case, it is beneficial to provide the local search with an alternative initial ordered list, which heuristic maps to a feasible schedule immediately.
4.2.1 Feasible Schedule for a Single Resource
We use the bin-packing CP formulation for monoprocessor (as described in [12]) relying on the height-divisible packing of Section 2.3.2. The CP solver processes one resource at a time and creates a feasible schedule for each one of them. Afterward, these feasible schedules are used to infer a suitable ordered list for the local search.
Set of tasks $\mathcal{J}_{m}$ (with periods $\tau(1),\dots,\tau(j),\dots,\tau(\rho)$) induces an associated set of rectangles $\mathcal{R}_{m}$, whose heights and widths are given as in Section 2.3.2. The set of rectangles’ heights corresponds to the periods of the tasks: $h_{j}=\tau(\rho)/\tau(j)$. Thus, $h_{1}=H$, which is equal to the height of the entire bin, and $h_{\rho}=1$.
Figure 5: Canonical form HD2D packing obtained by CP warm start. Bold T-like lines show how the bin is recursively partitioned into sub-bins.
Our formulation packs the rectangles of $\mathcal{R}_{m}$ in a height-divisible manner to produce a packing in canonical form; for each rectangle it holds, that all rectangles to its left have a height greater or equal. See the example in Fig. 5, which shows canonical packing corresponding to Fig. 2(b) (if there exists a feasible packing, then there exists a feasible canonical packing - see the proof in [12]). The T-like bold lines in the figure show how the bin is recursively partitioned into sub-bins, where rectangles with smaller heights can be packed, given the height-divisible property of packing. Rectangles with height $h_{1}=H=6$ can only be packed into a single sub-bin that spans the entire figure. The rectangles with height $h_{2}=3$ have two available sub-bins. Finally, the rectangles with $h_{3}=1$ can be packed into six different sub-bins. However, since only three of these sub-bin of height 1 are filled with rectangles, the remaining three have width 0 and are not shown in Fig. 5.
To utilize this sub-bin view of the packing, we partition the rectangles to lists $\mathcal{R}_{m}^{j}$ according to their height $h_{j}$ (an example is described later in the section). Each such list is associated with a vector of widths of the rectangles $\mathbf{l}^{j}$ and a vector of integer assignment variables $\mathbf{a}^{{j}}$ with domain $\{0,\cdots,(H/h_{j})-1\}$. The assignment variable determines which sub-bin the rectangle will be packed into. Furthermore, we associate each available sub-bin $k$ of height $h_{j}$ with one variable in the load vector $\mathbf{L}^{j}=(L^{j}_{0},\dots,L^{j}_{k},\dots,L^{j}_{(H/h_{j})-1})$. The load variable keeps track of how much the sub-bin is filled. The width of the bin $w$, into which we pack the rectangles, is given by the shortest period $\tau(1)$.
The height-divisible bin-packing problem is modeled as follows:
$\mathtt{pack}(\mathbf{L}^{j},\mathbf{a}^{j},\mathbf{l}^{j})$ $\forall j\in\{1,\dots,\rho\}$ (3)
$\sum_{j=1}^{\rho}L^{j}_{\lfloor k/h_{j}\rfloor}\leq w$ $\forall k\in\{0,\dots,H-1\}$ (4)
The pack constraint forces the model to assign each rectangle to one of the appropriate sub-bins and sets the associate load variable of the sub-bin $L^{j}_{k}$ to the sum of the widths of assigned rectangles.
For example, Fig. 5 contains five rectangles of height $h_{2}=3$ in $\mathcal{R}_{m}^{2}=(3,4,5,6,7)$ with corresponding widths given by $\mathbf{l}^{2}=(17,9,6,5,2)$. The top sub-bin for $h_{2}=3$ is occupied by rectangles 3, 4, and 7, resulting in $L^{2}_{1}=28$. The bottom sub-bin for $h_{2}=3$ is occupied by rectangles 5 and 6, resulting in $L^{2}_{0}=11$. Therefore, the vector of assignment variables is $\mathbf{a}^{{2}}=(1,1,0,0,1)$.
The second constraint enforces that for each row of height 1, the sum of loads of sub-bins that cross the row (across the different heights) does not exceed the bin’s width. In case of row 2 in Fig. 5, we have:
$L^{1}_{0}+L^{2}_{0}+L^{3}_{2}=12+11+13=36\leq 40=w$
Once the CP solver finds a feasible HD2D packing in canonical form (given the values of assignment variables $\mathbf{a}^{j}$), we transform it to the schedule of the original single resource PSP. The coordinates of the rectangles determine the start times of the associated tasks.
4.2.2 Inferring Ordered List from Schedule
The schedule provided by the CP is used to create an initial ordered list such that its processing by the FFS heuristics leads to a feasible solution (such a solution is always found for 100% utilized instances, but it is not always found for less utilized instances as shown in Section 4.2.3).
An ordered list that should produce a feasible solution for leftmost heuristic is obtained simply by sorting the tasks by their start time. However, to be used with the predecessor method, this relatively arbitrary ordered list may lead to unexpected behavior since it does not respect inter-resource coupling via precedences. We attempt to resolve this by employing an order-determining CP model for the predecessor method. In this model, each task $J_{i}^{k}$ of the instance is associated with integer variable $a^{k}_{i}$. All variables $a^{k}_{i}$ are members of set $\mathbf{a}$. The CP model is:
$\mathtt{AllDiff}(\mathbf{a})$ (5)
$a^{k}_{i}<a^{l}_{j}$ $\forall\mathcal{J}_{m}\leavevmode\nobreak\ \forall J_{i}^{k},J_{j}^{l}\in\mathcal{J}_{m}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }t_{i}^{k}<t_{j}^{l}$ (6)
$a^{k}_{i+1}<a^{k}_{i}$ $\forall\mathcal{C}^{k}\leavevmode\nobreak\ \forall J_{i}^{k},J_{i+1}^{k}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }t_{i}^{k}+p_{i}^{k}>t_{i+1}^{k}$ (7)
This model processes all the tasks (and thus also the resources) simultaneously. It directly determines the ordered list of tasks with the ”all-different” constraint - the value of $a_{i}^{k}$ is the position of $J_{i}^{k}$ in the ordered list. Furthermore, when the successor task $J_{i+1}^{k}$ would be scheduled later than planned due to the start time of its predecessor, we ensure it is processed before its predecessor so the heuristic does not unexpectedly postpone it.
4.2.3 Feasibility Issues
The aforementioned CP approach consistently finds feasible solutions for 100% utilized instances. Nevertheless, when the utilization drops, it may happen that the heuristic cannot transform the provided order into a feasible solution. However, this happens at worst for 10% of instances on the most affected instance set. Furthermore, as we show in Section 5, the local search successfully finds feasible solutions for less-utilized instances even without a warm start.
(a) Schedule obtained by CP solver.
(b) Partial infeasible schedule produced by the FFS heuristic for a given order of tasks deduced from CP schedule.
Figure 6: Schedules documenting the issue with a warm start on less-utilized resources. Notice the gap caused by scheduling task 123 at an incorrect time caused by FFS, leading to infeasibility.
This is demonstrated in Fig. 6. In Fig. 6(a), the schedule of one specific resource produced by the CP warm start is clearly feasible. However, when the FFS heuristic processes tasks row by row, we can see in Fig. 6(b) that it misplaces task 123; this CP model would put the task into the second row, but the heuristic schedules it in the first one, as it operates in a first fit manner, and there is excessive space available. This creates a gap due to the task’s periodicity. Finally, this inefficiency makes it impossible to schedule task 187 with a long processing time later.
4.3 Constraint Programming Matheuristic
The local search performs generally well, but it sometimes struggles to minimize the degeneracy of highly-utilized instances. We demonstrate this using Fig. 7 evaluating the local search on a set of more difficult instances. In Fig. 7(a), each curve shows the degeneracy of the best-known solution of instance as the local search progresses. We see that local search did not converge for several instances. However, in other cases, the local search rapidly decreased the degeneracy and stagnated for the rest of the experiment.
We investigated how the degeneracy is distributed within these stagnating instances. Let $\mathbf{J}_{m}^{\tau(\rho)}$ be set of tasks of machine $m$ with period $\tau(\rho)$. Similarly, let $\mathbf{C}^{\tau(\rho)}$ be sets of chains with period $\tau(\rho)$. Blue dots in Fig. 7(b) show the values of $\phi^{\tau(\rho)}=\frac{\sum_{\mathcal{C}^{k}\in\mathbf{C}^{\tau(\rho)}}D^{k}}{D_{sum}}[\%]$ for each such instance; instances are sorted in ascending order of $\phi^{\tau(\rho)}$ in the chart. The metric represents the percentage of the degeneracy left in the chains with the longest period (i.e., $T^{k}=\tau(\rho)$, the hyperperiod). To our surprise, these longest-period chains produce 40%-80% of degeneracy, even though they have just a single occurrence in the schedule.
Furthermore, by orange dots we show values of $\Omega^{\tau(\rho)}=\frac{\sum_{J_{i}^{k}\in\mathbf{J}_{m}^{\tau(\rho)}}p_{i}^{k}/T^{k}}{U_{m}}[\%]$; it shows the ratio of the utilization caused by the longest period tasks, calculated the most utilized resource $m$. We can see that this value is lower than $\phi^{\tau(\rho)}$. This suggests that while the longest-period tasks do not necessarily utilize the instance the most, their chains significantly contribute to the overall degeneracy. The culprit is probably the fact that the ordered list generated by the local search is generally almost rate-monotonic. Since the FFS heuristic processes these tasks later, there are fewer options to schedule them with low degeneracy.
(a) Degeneracy of the current best solution over time. Each line corresponds to a single instance.
(b) Ratio of degeneracy $\phi^{\tau(\rho)}$ and ratio of utilization $\Omega^{\tau(\rho)}$ covered by the longest-period chains and tasks.
Figure 7: Characteristics of instances not solved to optimality by local search.
This observation motivated us to use a matheuristic to improve the schedule further. Similarly, as in [36, 37], we focus on a natural way to isolate part of the problem: individual periods. If we fixed all tasks with periods shorter than $\tau(\rho)$, we could try to re-schedule the longest-period tasks within the complete hyperperiod $\left[0;\tau(\rho)-1\right]$. Other tasks behave as fixed obstacles. Similarly, we can focus on optimizing chains with arbitrary period $\tau(\beta)$. With this focus on a single period, we try to optimize the schedule using a CP solver, exploiting its capabilities for solving common scheduling problems using interval variables.
Let $\mathcal{J}_{m}^{O}$ be a set of tasks for the period $\tau(\beta)$ we are trying to optimize, and let $\mathcal{J}_{m}^{F}$ be a set of the tasks of other periods with their start time fixed. This is done simultaneously for all resources together but focusing on a single period only. As in Section 2.3, we shift each resource’s start time so that a task with the shortest period starts at time 0. Let $o_{m}$ be such offset. Again, let $s_{i}^{k}$ be the shifted core schedule of the task, as it is in Section 2.3. Finally, set of chains with period $T^{k}=\tau(\beta)$ is denoted as $\mathbf{C}^{\tau(\beta)}$.
We formulate the following matheuristic CP model. The model tries to re-schedule tasks with period $\tau(\beta)$ to decrease the criterion while keeping the other tasks fixed:
$\min\leavevmode\nobreak\ \sum_{\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}}(\lceil S^{k}/\tau(\beta)\rceil-1)$ (8)
$\mathtt{NoOverlap}(\mathcal{J}_{m}^{O}\cup\mathcal{J}_{m}^{F})$ $\forall{m}$ (9)
$\mathtt{StartOf}(J_{i}^{k})=s_{i}^{k}$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{F}$ (10)
$\mathtt{StartOf}(J_{i}^{k})\geq 0$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (11)
$\mathtt{EndOf}(J_{i}^{k})\leq\tau(\beta)$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{O}$ (12)
$ps_{i}^{k}=(\mathtt{StartOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (13)
$pe_{i}^{k}=(\mathtt{EndOf}(J_{i}^{k})+o_{\pi_{i}^{k}})\mod\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (14)
${df}_{0}^{k}=0$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (15)
$ps_{i+1}^{k}-pe_{i}^{k}\geq 0\implies{df}_{i+1}^{k}=ps_{i+1}^{k}-pe_{i}^{k}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k},J_{i+1}^{k}\in\mathcal{C}^{k}$ (16)
$ps_{i+1}^{k}-pe_{i}^{k}<0\implies{df}_{i+1}^{k}=ps_{i+1}^{k}-pe_{i}^{k}+\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k},J_{i+1}^{k}\in\mathcal{C}^{k}$ (17)
$S^{k}=\sum_{\forall J_{i}^{k}\in\mathcal{C}^{k}}p_{i}^{k}+{df}_{i}^{k}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (18)
$J_{i}^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{intervalVar}$ $\forall{m}\forall J_{i}^{k}\in\mathcal{J}_{m}^{F}\cup\mathcal{J}_{m}^{O}$ (19)
$S^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{integerVar}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (20)
$ps_{i}^{k},pe_{i}^{k},df_{i}^{k}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathtt{integerVar}$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}\forall J_{i}^{k}\in\mathcal{C}^{k}$ (21)
Eq. 9 ensures the non-overlapping of intervals on each resource. Eqs. 10, 11 and 12 either fix or limit the intervals of the tasks. Variables $ps_{i}^{k}$ and $pe_{i}^{k}$ correspond to the start- and completion time of the task once the resource offset is considered. We used them to calculate the consecutive differences within the chain $df_{i}^{k}$. This, together with tasks’ processing times, forms the start-to-end latency $S^{k}$ in equation Eq. 18.
Criterion precisely models the degeneracy $D_{sum}$. Note that both $D_{\max},D_{sum}({\alpha})$ can be modeled in a similar manner. Furthermore, we add the following constraint for sum-based criteria to ensure that the degeneracy of individual chains is not worsened, which we observed to improve the performance. Let $deg^{k}$ be the degeneracy of chain $\mathcal{C}^{k}$ given the current solution. Then we add the following constraints:
$S^{k}\leq(deg^{k}+1)\cdot\tau(\beta)$ $\forall\mathcal{C}^{k}\in\mathbf{C}^{\tau(\beta)}$ (22)
For large instances, the performance of the CP solver starts to diminish due to the large number of task intervals. Thus, when the size of the problem exceeds a defined bound, we do not optimize all the chains of given period $\tau(\beta)$ simultaneously; rather, we optimize only $K$ chains with the largest value of degeneracy (keeping others fixed). We repeat this with $K$ most problematic chains until a time limit is reached.
5 Experiments
We used C++ and Python 3.10. Experiments were performed on Intel Xeon E5-2690 using a single thread each time. CP Optimizer v22.1 was used as a CP solver.
The experiments were primarily performed on synthetically generated instances inspired by those in [30]. Initially, each resource is split into segments that become tasks of period $\tau(1)$ and segments that are recursively split further to allocate longer-period tasks. Base period $\tau(1)$ was set to 100, 200, or 400, and there were between 3 to 5 distinct periods per instance, with ratios of 2, 3, and 4 between consecutive periods. Chains are generated afterward, ensuring that the degeneracy of the instance is zero.
(a) Star topology
(b) Triangle topology
(c) Bridge topology
(d) Stubbed line topology
Figure 8: Topologies used for TOP instance sets.
Three kinds of instance sets were produced. Set GEN-x contains the instances for the general case, where the chain can pass from one arbitrary resource to another, even returning to the original one; processing times among tasks of the chain can differ, and x refers to $\min_{m}U_{m}$; the minimum utilization among the resources.
Sets TOP were generated using underlying topologies: star, triangle, bridge, and stubbed line (Fig. 8). TOP-L contains instances using the stubbed line topology; the rest is put into the set TOP-R. Processing time is the same for all tasks of each chain. The links are full-duplex, and messages are sent between endpoints. However, there are also tasks being sent between two neighboring nodes only to increase the utilization (when we omit these additional tasks in sets TOP-R, TOP-L, we obtain less-utilized sets TOP-R-L, TOP-L-L).
Finally, THEORY set is used to discuss our complexity results; it consists of line and out-tree topologies, where all chains start in the same common resource (see Section 3.2). The parameters of the sets are presented in Table 1.
# instances # resources avg # tasks avg # chains avg chain length avg $\min_{m}U_{m}$ GEN-1 1194 5-10 2193 254 8.4 1.00 GEN-0.98 1194 5-10 1786 208 8.3 0.98 GEN-0.96 1194 5-10 1558 183 8.2 0.96 GEN-0.94 1194 5-10 1374 163 8.1 0.94 GEN-0.9 1194 5-10 1093 131 7.9 0.90 TOP-L 796 46 7634 2247 2.7 1.00 TOP-L-L 796 46 6155 2246 2.7 0.71 TOP-R 2388 10-24 2487 792 2.3 1.00 TOP-R-L 2388 10-24 1929 792 2.35 0.73 THEORY 320 10-77 1102 96 11.2 0.66
Table 1: Parameters of synthetically generated instance sets.
5.1 Finding Feasible Solutions
First, we studied the effect of a warm start and the ability of the local search to find feasible and high-quality solutions. We ended with a strategy that uses the CP to warm-start the heuristic for instances with utilization close to 100%; otherwise, it relies on the local search itself. This strategy provided a good trade-off between the feasibility success rate and the value of the criterion.
The experiment was performed on TOP and GEN instances with a time limit of 6 minutes. In Table 2, we report the percentage of instances where the feasible solution was found, the average value of degeneracy $D_{sum}$, and a percentage of instances with 0-$D_{sum}$ solutions (when the final solution has degeneracy equal to 0). The results are reported for both leftmost and predecessor heuristic methods, either being warm-started with CP or starting from the default ordered list. The last method predecessor-flow follows our outlined flow in Fig. 4). It starts a local search from the default ordered list, and if it does not find a feasible solution within 15 seconds (giving local search sufficient time to initialize), it runs a CP warm start instead.
leftmost leftmost-CP predecessor predecessor-CP predecessor-flow % feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ GEN-1 21.4 583.5 0.0 100.0 802.1 0.0 0.1 651.0 0.0 85.9 793.0 0.0 85.8 792.2 0.0 GEN-0.98 100.0 312.8 0.0 89.4 625.5 0.0 100.0 143.4 3.0 73.3 341.9 0.0 99.0 149.4 3.2 GEN-0.96 100.0 259.6 0.0 98.3 511.4 0.0 100.0 29.4 26.0 83.7 144.6 0.6 100.0 33.7 25.0 GEN-0.94 100.0 232.7 0.0 99.1 437.9 0.0 100.0 6.5 51.8 86.4 65.4 2.4 100.0 7.7 50.6 GEN-0.9 100.0 190.7 0.0 100.0 326.4 0.0 100.0 0.9 80.2 87.4 21.3 13.5 100.0 0.9 81.2 TOP-L 69.1 919.4 0.0 100.0 836.1 0.0 23.7 258.1 0.0 82.8 766.8 0.0 80.8 731.6 0.0 TOP-L-L 100.0 808.2 0.0 85.3 773.8 0.0 100.0 20.2 84.7 75.8 267.3 2.4 99.7 27.8 81.3 TOP-R 69.0 135.6 2.0 99.9 158.4 6.7 64.2 11.0 25.0 98.0 142.0 2.8 99.0 93.6 22.3 TOP-R-L 100.0 101.5 3.9 92.0 113.8 29.3 100.0 0.0 99.7 91.2 8.9 70.9 100.0 0.0 99.8
Table 2: Results of non-warm-started and warm-started local search methods.
We can see that the predecessor FFS heuristic achieves great results if the utilization is below 98%. It finds a feasible solution and achieves low values of $D_{sum}$. In case of predecessor-CP, CP solver finds a feasible schedule in most cases (99% for when solving one resource after another), but it may not produce an ordered list acceptable by the FFS heuristic (see Section 4.2.2). Even when the ordered list succeeds, final values of $D_{sum}$ seem to be worse than in the case of predecessor results. The reason may be that the initial ordered list lacks rate-monotonicity, making it harder for local search to improve it. However, CP warm start helps to solve the most utilized instances among our sets TOP-L and GEN-1.
The results of the hybrid predecessor-flow approach are provided in the last column of Table 2. We can see that even this simple strategy, applying CP warm start when needed (as outlined in Fig. 4), maintains high success rates and low degeneracy values. Thus, we use this approach in the rest of the paper.
Finally, we also tinkered with using the genetic algorithm to find a suitable ordered list of tasks. However, we did not pursue it further; the ordered lists obtained by the genetic algorithm’s operators too often produced infeasible solutions, and within the same number of evaluations, the genetic algorithm found either worse or no solution at all. Given the constrained characteristics of our problem and the use of the FFS heuristic, the local search appears to be much more efficient.
5.2 Benefit of Matheuristic
In this section, we discuss the benefits of the matheuristic CP, which is used to improve the degeneracy of the feasible schedule found by local search. Firstly, we determined that for a larger number of tasks, the best course of action is to optimize only several chains of the same period at once. Empirically, we set the number of non-fixed tasks of the model to at most 500 and the time limit for each optimization iteration to 10 seconds.
We compared two ways how to use matheuristic: (i) fixed matheuristic, when the matheuristic was initiated after five minutes of local search; and (ii) dynamic matheuristic, when the matheuristic starts once the local search does not improve the solution for 60 seconds. We determined this dynamic threshold from the small experiment, studying how the local search improves the solution over time (see Fig. 7(a)). The local search with no matheuristic is used as a baseline. In all cases, the computation time was six minutes.
In Table 3, we compare the degeneracy obtained while using fixed or dynamic matheuristic ($D_{sum}^{\mathrm{MH}}$) and the degeneracy obtained by the local search alone ($D_{sum}$). We report averaged values of $D_{sum}^{\mathrm{MH}}$ and differences $\Delta D_{sum}=D_{sum}^{\mathrm{MH}}-D_{sum}$. Thus, a negative value of $\Delta D_{sum}$ means that, on average, using matheuristic was beneficial.
fixed matheuristic dynamic matheuristic avg $D_{sum}^{\mathrm{MH}}$ avg $\Delta D_{sum}$ avg $D_{sum}^{\mathrm{MH}}$ avg $\Delta D_{sum}$ GEN-1 642.0 -151.0 687.0 -106.0 GEN-0.98 128.1 -15.3 130.2 -12.6 GEN-0.96 32.3 2.9 27.1 -2.3 GEN-0.94 6.4 -0.1 6.4 -0.1 GEN-0.9 0.9 0.0 1.0 0.1 TOP-L 585.8 -180.2 684.4 -45.2 TOP-L-L 23.8 3.6 25.5 -2.3 TOP-R 55.8 -86.6 63.2 -29.7 TOP-R-L 0.0 0.0 0.0 0.0
Table 3: Effect of fixed and dynamic matheuristic on values of $D_{sum}$.
Especially for the maximally-utilized instances, the benefit of the matheuristic is significant - the newly obtained values of $D_{sum}^{\mathrm{MH}}$ are significantly lower than the previous values $D_{sum}$. However, for the fixed matheuristic, its benefit drops significantly and even diminishes for some instances - the local search probably has not yet converged, and running it further could yield a better solution. This can be seen in Fig. 9. The histogram of values $\Delta D_{sum}$ is shown for instances GEN-0.96 (for clarity, only those with $\Delta D_{sum}\neq 0$ are presented). We can see that the majority of the differences are negative, meaning the matheuristic was beneficial. However, a few instances with very large positive values of $\Delta D_{sum}$ skew the average to the reported value of $2.9$.
Figure 9: Histogram of values of $\Delta D_{sum}$ for instance set GEN-0.96 and fixed matheuristic.
While the dynamic setting does not produce as good results as the fixed one for the most highly-utilized instances, we observed that its results are more consistent across the instance sets; e.g., on GEN-0.98, the worst result reported by the fixed setting was the value of $D_{sum}$ greater by 253 than the one reported by the local search only. In contrast, the worst difference reported by dynamic matheuristic was only 19. Therefore, we used the dynamic matheuristic in the rest of the paper.
5.3 Time Complexity of Special Cases
In this section, we looked into scheduling for line and out-tree topologies. We illustrate how well the offset schedule approach of Section 3.2 outperforms our proposed methodology in these special cases.
To do this, we utilized the set THEORY with 320 instances, which upholds the properties outlined Section 3.2. We compare the best results obtained by our proposed flow of Section 4 with the offset schedule approach described in Section 3.2.1: we attempt to find a feasible schedule using the CP model of Section 4.2 for the first common resource among the chains, and then we copy the schedule to other resources with appropriate offsets.
predecessor-flow offset schedule % feasible avg $D_{sum}$ 0-$D_{sum}$ feasible avg $D_{sum}$ 0-$D_{sum}$ THEORY 77.8 236.2 0.0 90.1 0 90.1
Table 4: Comparison of feasible solutions found and final degeneracy obtained by predecessor-flow and offset schedule on THEORY instances.
In Table 4, we can see that the offset schedule wins in this special case, finding more feasible solutions, each with zero degeneracy. The reason is that it only needs to find a feasible solution for a single resource. Then, it immediately obtains a solution with optimal degeneracy. This highlights the crucial role of problem analysis, which can massively simplify PSP.
5.4 Different Criteria
Up until now, we have focused mostly on degeneracy $D_{sum}$ as the criterion. In this section, we briefly study the differences in results when we optimize $D_{sum}({0.75})$ or $D_{\max}$ instead. We ran the local search for instances of set GEN-O.98 with the two aforementioned criteria and compared the results with the baseline results, optimizing $D_{sum}$.
(a) Minimizing $D_{sum}$
(b) Minimizing $D_{sum}({0.75})$
(c) Minimizing $D_{\max}$
Figure 10: Values of metrics for the instances of GEN-0.98 for different minimization criterion. Instances were sorted in ascending order of $D_{sum}$.
Results per instance are shown in Fig. 10. In each figure, three vertically aligned markers show the values of the metrics for a single instance. Chart (a) shows the values of metrics when $D_{sum}$ was minimized, chart (b) corresponds to the minimization of $D_{sum}({0.75})$, and chart (c) to the minimization of $D_{\max}$.
From all the figures, it is clear that values of $D_{sum}({0.75})$ are highly correlated with the original degeneracy in all experiments, even though the used instances do not necessarily admit solution with $D_{sum}({0.75})=0$ (but they do admit solution with $D_{sum}=0$). The results obtained by minimizing $D_{sum}$ in Fig. 10(a) and $D_{sum}({0.75})$ in Fig. 10(b) are rather similar. However, when $D_{\max}$ is used as an optimization criterion in Fig. 10(c), the worst reported value of $D_{\max}$ among the instances is 8 instead of 9.
Averaged values of the metrics are shown in Table 5. There, the difference between optimizing $D_{sum}$ and $D_{sum}({0.75})$ is clear. Interestingly, direct optimization of $D_{sum}({0.75})$ actually led to an overall worse solution with respect to all the metrics than when we used the $D_{sum}$ criterion, even though the space of the feasible solutions is the same. One of the reasons for this can be the local search procedure; the initial phase, where the chains are reordered, ends after the first failure to improve the criterion. This happens earlier in the case of the $D_{sum}({0.75})$. We believe there might be a connection between the worse performance of optimization with stricter $D_{sum}({\alpha})$ criterion and the difficulty of finding feasible solutions in a TSN setting with release times and deadlines.
criterion metric $D_{sum}$ $D_{sum}({0.75})$ $D_{\max}$ avg $D_{sum}$ 143.4 175.6 206.5 avg $D_{sum}({0.75})$ 274.9 296.7 347.8 avg $D_{\max}$ 5.07 5.35 4.70
Table 5: Average values of metrics given the optimization criterion. Results reported for instance set GEN-0.98.
5.5 Real Life-like Instances
Finally, we performed several experiments on real-life inspired instances, which highlight how well our approach performs in setting much closer to the practical applications.
5.5.1 Production Line
Firstly, we used production-line instances from [29]. In that paper, three types of instances were generated, one of which used harmonic periods. The original paper handled problems with more complex instances than we do, including frame-isolation constraints, queues, release times, and deadlines, and reported low success rates once the maximum utilization among the resources rose above 50%. These additional constraints are omitted for the purpose of this paper.
% feasible avg $D_{sum}$ 0-$D_{sum}$ production-line 100.0 530.9 80.5
Table 6: Feasibility and degeneracy values reported for production-line dataset (3000 instances) optimized with predecessor-flow.
In Table 6, we present the results across the 3000 mentioned instances, obtained after six minutes. For most instances, we find a 0-$D_{sum}$ solution. For the rest of the instances, the values of degeneracy were quite high. However, in these cases, the number of tasks was, on average, 30000, and the number of chains was 2700. The average $D_{\max}$ was equal to $6.2$, i.e., the most delayed chain reported degeneracy $<7$. Altogether, better results could be achieved by providing the solver more time, as in these cases, optimization has not yet converged. However, for these large-scale instances, it would probably be worth investigating more aggressive neighborhood operators or decomposition techniques. Note that CP warm start was still able to determine feasible schedules even for these large instances.
5.5.2 Avionic
The second real-life set is the avionic instances, also from [29]. Chains of the five industrial instances from the avionic domain were scheduled across 70-92 resources, with a maximum utilization of less than 38%. Thus, we also created a second set of instances in which the periods were halved, doubling the utilization. However, in both cases, our approach found a feasible 0-$D_{sum}$ solution within two minutes (given the constraints we study in this paper).
5.5.3 Automotive
Finally, we generated a set of automotive instances following a zonal architecture in modern cars. An example of such architecture is shown in Fig. 11(a), which we used as an inspiration for our network in Fig. 11(b). The central controller used 5Gb/s communication speed, the encompassing ring used 2.5Gb/s, and the rest of the connections used 1 Gb/s. We generated messages between leaf and central nodes of a size from 72 to 1542 bytes, as is used in Ethernet frames. The messages’ periods were 10 ms, 50 ms, 250 ms, and 500 ms.
(a) Example of zonal architecture [38].
(b) Zonal-architecture inspired network used in the experiment.
Figure 11: Example of zonal architecture and used network.
We generated 20 such instances with an average maximum utilization of $U_{m}=0.9$ and 300000+ tasks each. However, due to the maximum length of the Ethernet messages, their processing times were short in contrast to the minimum period. Thus, our approach found feasible solutions with $D_{\max}\leq 2$ within 6 minutes, even for these large-scale instances. Even though it managed to perform only a few iterations of the local search, it managed to produce a schedule with satisfactory characteristics.
6 Conclusion
In this paper, we presented complexity results related to the periodic scheduling problem with harmonic periods and dedicated resources. Crucially, we have shown that for communication over the line and tree topologies with a single source node, an offset schedule approach can be used to efficiently find a high-quality solution.
We proposed a warm start approach, which successfully helps the local search optimize instances with very high utilization. Furthermore, we described a CP matheuristic, which leverages CP scheduling capabilities by optimizing the schedule while targeting one period at a time.
Several interesting conclusions can be drawn from our results. Even though scheduling periodic tasks is the principal component of TSN, we achieved much higher utilization rates in our restricted problem than is typical in TSN applications; it seems that the additional constraints introduced by the TSN standard are the main culprit behind the difficulty of scheduling, not just the PSP alone.
Another interesting property is the relationship between the number of tasks and the difficulty of solving the instance. We had problems even generating instances that were not quickly solved by the baseline heuristic. For a fixed size of the ”observational interval”, increasing the number of tasks will, at one point, necessarily decrease the average processing time of tasks, which generally makes them easier to schedule.
Finally, we successfully demonstrated how understanding the instances’ characteristics can be exploited in the PSP domain. To name a few, using specialized methods for specific resource topologies or focusing on optimizing an isolated sub-problem, which was, in our case, a set of tasks corresponding to a single period.
Acknowledgments
This work was supported by the Grant Agency of the Czech Republic under the project GACR 22-31670S and by the Grant Agency of the Czech Technical University in Prague, grant No. SGS22/167/OHK3/3T/13.
This work was co-funded by the European Union under the project ROBOPROX (reg. no. CZ.02.01.01/00/22_008/0004590).